rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#12396 — FS#17022 — nwk <> ldn

Attached to Project— Network
Incident
Whole Network
In progress
0%
We currently have 1 link down between pob-1-n7 and nwk-1-n7.

We have isolated the port-channel in question and it will remain this way until the problem is fixed.

Our supplier is referring a fiber cut:

03/11/2016 18:33:24 GMT - The Transport NOC has advised that a suspected fiber cut in New York, NY is impacting unprotected transport services. Field Services have been engaged to test fiber to locate the fault point. There is no ETR available at this time. Please be advised that updates for this event will be relayed on an hourly basis unless otherwise noted.
Comment by OVH - Monday, 14 March 2016, 17:06PM

last update :

03/11/2016 11:09 PM GMT

*** CASCADED EXTERNAL NOTES 11-Mar-2016 23:07:05 GMT From CASE: 10517327 - Event
The Local Provider has reported that the city authorities are still working on the manhole and inspecting to ensure the site is safe for teams, before access will be granted. The Local Provider has a team on site and replacement 432-count fiber cable is currently en route. Once site access is granted, and the fiber cable arrives, the Local Provider will assess the damage to determine an appropriate repair plan and commence cable prep work. There remains no ETR available.

03/11/2016 9:39 PM GMT

*** CASCADED EXTERNAL NOTES 11-Mar-2016 21:37:17 GMT From CASE: 10517327 - Event
The Local Provider has advised that local authorities are not allowing any access to the manhole at this time. Once the local authorities have addressed the trouble in the manhole and made the site safe for entry, the Local Provider will be granted access. Due to the severity of the trouble, and the jurisdiction of city authorities, there is no ETR available. Please be advised that further updates will be provided as new information becomes available.

03/11/2016 9:13 PM GMT

*** CASCADED EXTERNAL NOTES 11-Mar-2016 21:03:41 GMT From CASE: 10517327 - Event
The Local Provider has advised that the failure is likely due to a manhole fire. The Local Provider's field teams are investigating at this time and there remains no ETR available. New information will continue to be relayed as it becomes available.

03/11/2016 9:10 PM GMT

*** CASCADED EXTERNAL NOTES 11-Mar-2016 20:46:15 GMT From CASE: 10517327 - Event
The NOC continues to chase the Local Provider for the details and status of the fiber issue on their network. No ETR has been provided. Please be advised that new information will be relayed as soon as it becomes available.

03/11/2016 8:46 PM GMT

*** CASCADED EXTERNAL NOTES 11-Mar-2016 20:17:06 GMT From CASE: 10517327 - Event
Field Services tested the fiber and results indicate a failure on the Local Providers network. The Local Provider has been engaged for investigation and status of the trouble. There is no ETR available at this time.


Comment by OVH - Monday, 14 March 2016, 17:07PM

The local provider reported a 700 foot section of 432 count cable has been put into place to bypass the fibers damaged by fire. Local provider field technicians continue preparations to begin splicing with a total ETTR of 19:00 GMT. Updates will be sent as information is received.


Comment by OVH - Monday, 14 March 2016, 17:07PM

The local provider reported fiber restorations efforts continue with an ETTR that remains unchanged, 19:00 GMT.


Comment by OVH - Monday, 14 March 2016, 17:08PM

Repairs are progressing at the site with the estimated completion time of 19:00 GMT.


Comment by OVH - Monday, 14 March 2016, 17:12PM

Our 10G are up again, we are reactivating the traffic but we continue to monitor closely before they have confirmation of the end of the incident.
——
The Local Provider has stated that fiber repairs are nearing completion but still ongoing. Please be advised that this event remains an active outage and updates will continue to be provided until all alarms are clear and the Local Provider's repair work is completed.